A discrete random variable π‘Œ has a cumulative probability distribution as shown below.


π‘Œ = 𝑦 0 1 2 3

𝑃(π‘Œ ≀ 𝑦) 5π‘˜ 10π‘˜ 11π‘˜ 12π‘˜


(a) Determine the value of π‘˜.

(b) Calculate expected value of π‘Œ.

(c) Calculate variance and standard deviation of π‘Œ.


1
Expert's answer
2022-02-22T13:47:09-0500

(a) Find the value of k from the condition

"\sum {{p_i}} = 1"

Then

"5k + 10k + 11k + 12k = 1 \Rightarrow 38k = 1 \Rightarrow k = \frac{1}{{38}}"

The distribution series has the form

"\begin{matrix}
y&0&1&2&3\\
p&{\frac{5}{{38}}}&{\frac{{10}}{{38}}}&{\frac{{11}}{{38}}}&{\frac{{12}}{{38}}}
\end{matrix}"

Answer: "k = \frac{1}{{38}}"

(b) "E(Y) = \sum {{y_i}} {p_i} = \frac{{0 \cdot 5 + 1 \cdot 10 + 2 \cdot 11 + 3 \cdot 12}}{{38}} = \frac{{68}}{{38}} = \frac{{34}}{{19}}"

Answer: "E(Y) = \frac{{34}}{{19}}"

(c) Let's find the variance:

"D(Y) = M({Y^2}) - {M^2}\left( Y \right) = \frac{{0 \cdot 5 + 1 \cdot 10 + 4 \cdot 11 + 9 \cdot 12}}{{38}} - {\left( {\frac{{34}}{{19}}} \right)^2} = \frac{{383}}{{361}}"

Then standard deviation is

"\sigma \left( Y \right) = \sqrt {D(Y)} = \sqrt {\frac{{383}}{{361}}} = \frac{{\sqrt {383} }}{{19}}"

Answer: "D(Y) = \frac{{383}}{{361}}" ; "\sigma \left( Y \right) = \frac{{\sqrt {383} }}{{19}}"